OpenStudy (anonymous):

A tortoise is walking in the desert. It walks for 48.8 meters at a speed of 3 meters per minute. For how many minutes does it walk?

OpenStudy (abhisar):

Time = \(\huge\frac{Distance}{Speed}\)

OpenStudy (abhisar):

Time = 48.8/3 =16.26 minutes

OpenStudy (abhisar):

got it ?

OpenStudy (abhisar):

You probably didn't notice, but you posted your \(\large\sf\color{blue}{\underline{\href{ /study#/groups/mathematics}{Mathematics }}}\) question in the Biology section. Next time when you ask a question, check if you're in the right section :)